vinegar contains 5.00% acetic acid by mass and has a density of 1.01g/ml. What mass(in grams) of acetic acid is present in 5.00 L of vinegar?
1
Expert's answer
2012-07-17T08:57:44-0400
The mass of vinegar is
m(vinegar) = V(vinegar) * d(vinegar) = 5 L * 1.01
g/mL = 5050 g
The mass of acetic acid is
m(CH3COOH) = w(CH3COOH) *
m(vinegar) = 0.05 * 5050 g = 252.5 g